OpenStudy (anonymous):

medal awarded In a group of 50 patrons, 8 patrons like lattes and espressos, 12 patrons like espressos and cappuccinos, 9 patrons like lattes and cappuccinos, and 5 patrons like all 3 coffee drinks. Altogether, 20 patrons like lattes, 28 patrons like espressos, and 19 patrons like cappuccinos. How many patrons don't like any of these coffee drinks? A. 9 B. 4 C. 3 D. 7

OpenStudy (anonymous):

Something doesn't add up. Total Patrons: 50 Lattes: 8 + 9 + 5 = 22 Espressos: 8 + 12 + 5 = 25 Cappuccinos: 12 + 9 + 5 = 26 What am I not seeing?

OpenStudy (anonymous):

i know right. this is very confusing

OpenStudy (anonymous):

Don't star my response I'm as lost as you are!

OpenStudy (anonymous):

:P ughhh this is horrible

OpenStudy (anonymous):

@driftracer305

OpenStudy (driftracer305):

woah......lol hot mess.....

OpenStudy (anonymous):

-singing- apex you suckkkkk really bad why to you make questions that making no scence at all why do u make my life a living helllllll. most of the the its like what the f**k is thissssss sh*ttttttttt

OpenStudy (driftracer305):

rofl.........

OpenStudy (anonymous):

yeah had to say that even though i messed up a bit